OpenStudy (anonymous):

My answer is B Solve 4x + 7 = 6x - 1 (2 points) Answer x ≈ 0.04 x ≈ 0.05 x ≈ 19.51 x ≈ 28.35

OpenStudy (anonymous):

6x-4x = 7+1 2x = 8 x = 4 all ur options are wrong

OpenStudy (anonymous):

either all your options are wrong or the question is wrong

OpenStudy (anonymous):

agree...acc to the question, i have the corret answre

OpenStudy (anonymous):

Solving Exponential Equations with Unequal Bases

OpenStudy (anonymous):

It's written like: Solve 4 ^ x + 7 = 6 ^ x – 1

OpenStudy (anonymous):

ok i tried to do this but still idk if something is wrong with it so eventually i substituted all the options into the equation and not one is correct